OpenStudy (lpruitt4):

Drew is selling candy bars for school. He started with 69 candy bars 2 weeks ago, and now he only has 26 left. How many candy bars did Drew sell? 43 95 42 26

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (raffle_snaffle):

so we started with 69 and we have 26 left. So between -, +, /, and * what do we use?

OpenStudy (raffle_snaffle):

let a = what he started with and let b = what he ended with b - a = ?

OpenStudy (raffle_snaffle):

b - a = what drew sold

OpenStudy (lpruitt4):

i think 69

OpenStudy (skullpatrol):

But he STARTED with 69?

OpenStudy (lpruitt4):

i think it is 26

OpenStudy (lpruitt4):

it is 23

OpenStudy (raffle_snaffle):

oops had that totally backwards. a - b = what he sold

OpenStudy (raffle_snaffle):

let a = what drew started with and let b = what he ended with ====> a - b = ? what drew sold

OpenStudy (skullpatrol):